-(3-8p)greater than or equal to 4p+25
Answer by checkley79(3341)

-(3-8p)>= (4p+25)
-3+8P>=4P+25
8P-4P>=25+3
4P>=28
P>=28/4
P>=7 ANS.
PRROOF:
LET P=7.01
-(3-8*7.01)>=(4*7.01+25)
-(3-56.08)>=28.04+25
-3+56.08>=53.04
53.08>=53.04